# How Real Are Synthetic Therapy Conversations? Evaluating Fidelity in Prolonged Exposure Dialogues

### v3

@@ Abstract
The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and high annotation costs. This work explores the use of synthetic Prolonged Exposure (PE) therapy conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models. We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity. We introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency. Our findings show that while synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it often struggles to capture the subtle dynamics of therapeutic interactions. Synthetic therapy dialogues closely match the structural features of real conversations (e.g., speaker switch ratio: 0.98 vs. 0.99), but often fails to adequately reflect key fidelity markers such as distress monitoring. This work highlights gaps in current evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures. Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
@@ 1 Introduction
Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022). Access to real clinical conversations—crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding—is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing. Synthetic data has emerged as a promising alternative, offering scalability and privacy preservation while reducing dependence on real-world annotations (Aher et al., 2023).
In trauma-focused mental health care, particularly Prolonged Exposure (PE) therapy for PTSD, large language models (LLMs) can generate synthetic therapy dialogues at scale. However, questions remain about whether these dialogues capture more than surface-level fluency—specifically, the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024). To address this, we develop and validate methods to measure clinically-relevant fidelity in generated dialogues—moving beyond standard metrics like coherence or perplexity. Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig. 1). Without clinical expertise, both humans and automated metrics tend to overestimate fidelity. We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS monitoring). This framework integrates automated scoring and expert clinical assessment for a more rigorous evaluation of synthetic dialogue quality. We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions—analyzing their linguistic, structural, and clinical fidelity characteristics. Our findings demonstrate the current strengths and limitations of synthetic dialogues and inform future improvements in generation and evaluation methods. Our study contributes to bridging the real–synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
@@ 3.1 Synthetic Dataset Generation and Evaluation
Recent work has explored synthetic clinical dialogue generation and assessment. BOLT evaluates LLM-generated therapist behaviors in general counseling (Chiu et al., 2024), while SimPsyDial benchmarks synthetic data using the Working Alliance Inventory (Qiu and Lan, 2024). CPsyCoun reconstructs dialogues for evaluation with BERTScore, GPTScore, and qualitative review (Zhang et al., 2024, 2019). Other studies focus on counselor style (Xie et al., 2024) and data augmentation (Kim et al., 2024). While Thousand Voices of Trauma introduces structural variation (BN et al., 2025), most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences. Our work directly addresses this by evaluating whether synthetic PE sessions capture clinical realism per PE protocols (e.g., SUDS, avoidance redirection), highlighting procedural errors missed by standard metrics.
@@ 3.2 Beyond Lexical Metrics
Evaluation is shifting from lexical metrics like BLEU (Papineni et al., 2002) to semantic ones such as BERTScore and GPTScore (Zhang et al., 2024; Xie et al., 2024). Psychological measures (CTRS, PANAS) in COUNSELINGEVAL assess empathy and coherence (Lee et al., 2024a), and interactional features like PQA are explored (Shen et al., 2024). However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW)—potentially useful for structured PE flows—remain underused. While some qualitative reviews consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and structured questioning is explored (Ren et al., 2024), robust frameworks for quantitative evaluation of PE-specific structure are lacking.
@@ 3.4 Bridging the Gap in PE Therapy Evaluation
PE is a first-line treatment for PTSD (Sherrill and Rauch, 2019; Varkovitzky et al., 2018; Rauch et al., 2021; Ragsdale et al., 2020; Evans et al., 2020; Yasinski et al., 2017). Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation, a gap remains in fidelity assessment for trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024). Current approaches lack metrics and focus tailored to PE’s unique conversational structure—including avoidance handling, trauma cue processing, and imaginal exposure flow—and do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024). Our work systematically evaluates synthetic data for alignment with PE structure and clinical validity, identifying areas for improving synthetic PE dialogue quality (see Table 1, Appendix A).
@@ 4 Methodology
We analyzed 400 prolonged exposure therapy conversations—200 real-world and 200 synthetic. The synthetic data comes from the publicly available Thousand Voices of Trauma dataset (BN et al., 2025), generated using Claude Sonnet 3.5 (Anthropic, 2024) with PE-specific prompting frameworks adapted from clinical guidelines. This data includes diverse therapist-client interactions across trauma types, therapy phases, and demographics.
The real-world PE sessions were collected under IRB-approved protocols, with participant consent, and cannot be released due to privacy constraints. Each session (1–1.5 hours) was transcribed using Amazon HealthScribe (Services, 2023), manually verified, and reformatted by merging consecutive speaker turns for readability. Each conversation followed a validated therapy fidelity checklist (See Table 1) to align with real-world standards. Both datasets underwent the same preprocessing:
@@ 4.1 Metrics and Analysis
To systematically compare real and synthetic therapeutic conversations, we selected a diverse set of linguistic, structural, and statistical metrics. These metrics provide insights into conversational dynamics, protocol adherence, and overall fidelity, ensuring a holistic evaluation of synthetic dialogue generation. To reduce subjectivity, we adopt a fidelity checklist (Table 1) and PE-specific metrics (Sec. 5) grounded in existing clinical guidelines for evaluating therapeutic adherence. Our methodology consists of four key analyses: system-level metrics comparison, correlation analysis, statistical significance testing, and PE-specific metrics.
@@ 4.1.1 System-Level Metrics Comparison
    By comparing means and standard deviations, we evaluate whether synthetic conversations approximate real dialogue structures. Significant deviations signal potential areas where synthetic data falls short.
@@ 4.1.2 Correlation Analysis
While mean comparisons provide a general overview, correlation analysis (see Table 3) quantifies the consistency of relationships across metrics between real and synthetic data. This determines if real-data patterns persist in synthetic dialogues. For instance:
  1. 1.
    Turn-taking consistency: Strong correlations in Normalized Therapist/Client Turns and Turn Ratios indicate that synthetic dialogues replicate real-world conversational structures.
  2. 2.
    Linguistic coherence: A high correlation in Utterance Length and Flow Entropy suggests that synthetic responses maintain natural variation in sentence length and randomness. Lower correlation values suggest that synthetic data may lack fidelity in preserving conversational nuances, particularly in aspects such as vocabulary richness and readability.
@@ 4.1.3 Statistical Significance Testing
To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 4). This assesses if variations are random or due to inherent inconsistencies. Key takeaways include:
  1. 1.
    Significant differences in linguistic complexity: Metrics such as Utterance Length, Turn Duration, and Vocabulary Richness show statistically significant differences, indicating that synthetic dialogues do not yet fully capture the expressive range of real conversations.
  2. 2.
    Minimal differences in turn-taking ratios: Therapist-Client Turn Ratio and Speaker Switches show weaker statistical significance, suggesting that structural patterns are relatively well-preserved.
  3. 3.
    Higher variability in generated text: The difference in Perplexity and Flow Entropy highlights challenges in maintaining linguistic variability without excessive repetition or over-simplification.
@@ 4.1.5 Qualitative Fidelity Assessment
To supplement automated evaluation metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using established PE clinical guidelines. Manual fidelity annotations were then reviewed by a licensed clinical psychotherapist for adherence to PE protocol. While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses—especially those potentially overlooked by automated metrics, rather than as quantitative ground truth. Representative examples of such lapses identified through this process are illustrated in Figure 1 and discussed in Section 6.2.
@@ 5.1 System-Level Metrics & Correlation
    Concise and structured responses: Synthetic dialogues are shorter (Average Utterance Length: 22.90 ± 1.74 vs. 68.72 ± 26.61) and more consistent (Utterance Length Std Dev: 18.54 ± 2.35 vs. 135.85 ± 66.25) due to LLM output constraints. Techniques like Chain-of-Thought prompting helps improve coherence, though larger output contexts are needed.
@@
    Correlation analysis: High correlations in turn-taking metrics (0.65–0.78 synthetic vs. 0.85–0.92 real) confirm accurate conversational dynamics. Lower correlations in linguistic complexity metrics (e.g., Readability Score, Vocabulary Richness, Perplexity) reflect structural differences but do not hinder the coherence/naturalness of the conversation.
@@
Norm. Total Turns $1.00\pm 0.0$ $1.00\pm 0.0$ $0.78$ $0.65$
@@
Turn Ratio (T/C) $0.01\pm 0.0$ $0.01\pm 0.0$ $0.80$ $0.66$
Word Ratio (T/C) $0.01\pm 0.0$ $0.00\pm 0$ 0.0 $0.82$ $0.68$
@@ 5.2 Statistical Significance Testing
T-C Turn Ratio $2.13\times 10^{3}$ $1.22\times 10^{-9}$
@@ 5.3 Feature Importance
Feature importance analysis identifies the metrics most influencing synthetic data generation, confirming trends from previous evaluations. While not the central focus, these results support the model’s alignment with conversational structure and linguistic patterns.
@@
    Lexical richness is moderately important: Vocabulary Richness plays a secondary role, reflecting that synthetic dialogues maintain varied but structured language. While real conversations are more flexible, synthetic responses balance vocabulary diversity with clarity.
@@ 5.4 Prolonged Exposure Specific Metrics
Table 5 presents the statistical test results comparing synthetic Prolonged Exposure data with real therapy session data. The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data. These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions. However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement. Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease. Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses. Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications. Addressing dynamic narrative evolution could further improve alignment with real data. However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics. For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring. These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.

### v4

@@ Abstract
Synthetic data adoption in healthcare is driven by privacy concerns, data access limitations, and high annotation costs. We explore synthetic Prolonged Exposure (PE) therapy conversations for PTSD as a scalable alternative for training clinical models. We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics like turn-taking and treatment fidelity. We introduce and evaluate PE-specific metrics, offering a novel framework for assessing clinical fidelity beyond surface fluency. Our findings show that while synthetic data successfully mitigates data scarcity and protects privacy, capturing the most subtle therapeutic dynamics remains a complex challenge. Synthetic dialogues successfully replicate key linguistic features of real conversations, for instance, achieving a similar Readability Score (89.2 vs. 88.1), while showing differences in some key fidelity markers like distress monitoring. This comparison highlights the need for fidelity-aware metrics that go beyond surface fluency to identify clinically significant nuances. Our model-agnostic framework is a critical tool for developers and clinicians to benchmark generative model fidelity before deployment in sensitive applications. Our findings help clarify where synthetic data can effectively complement real-world datasets, while also identifying areas for future refinement.
@@ 1 Introduction
Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022). Access to real clinical conversations, which are crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding, is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing. Synthetic data has emerged as a promising alternative, offering scalability and privacy preservation while reducing dependence on real-world annotations (Aher et al., 2023).
In trauma-focused mental health care, particularly Prolonged Exposure (PE) therapy for PTSD, large language models (LLMs) can generate synthetic therapy dialogues at scale. However, questions remain about whether these dialogues capture more than surface-level fluency, specifically the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024). To address this, we develop and validate methods to measure clinically relevant fidelity in generated dialogues, moving beyond standard metrics like coherence or perplexity. Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025b), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig. 1). Without clinical expertise, both humans and automated metrics tend to overestimate fidelity.
We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS (Subjective Units of Distress) monitoring). This framework integrates automated scoring and expert clinical assessment for a more rigorous evaluation of synthetic dialogue quality.
We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions, analyzing their linguistic, structural, and clinical fidelity characteristics. Our findings demonstrate the current strengths and limitations of synthetic dialogues and inform future improvements in generation and evaluation methods. Our study contributes to bridging the real-synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
@@ 3.1 Synthetic Dataset Generation and Evaluation
Recent work has explored synthetic clinical dialogue generation and assessment. BOLT evaluates LLM-generated therapist behaviors in general counseling (Chiu et al., 2024), while SimPsyDial benchmarks synthetic data using the Working Alliance Inventory (Qiu and Lan, 2024). CPsyCoun reconstructs dialogues for evaluation with BERTScore, GPTScore, and qualitative review (Zhang et al., 2024, 2019). Other studies focus on counselor style (Xie et al., 2024) and data augmentation (Kim et al., 2024). While The Pursuit of Empathy highlights the importance of empathy-aware evaluation (BN et al., 2025a), Thousand Voices of Trauma introduces structural variation (BN et al., 2025b), but most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences. Our work directly addresses this by evaluating whether synthetic PE sessions capture clinical realism per PE protocols (e.g., SUDS, avoidance redirection), highlighting procedural errors missed by standard metrics.
@@ 3.2 Beyond Lexical Metrics
Evaluation is shifting from lexical metrics like BLEU (Papineni et al., 2002) to semantic ones such as BERTScore and GPTScore (Zhang et al., 2024; Xie et al., 2024). Psychological measures (CTRS, PANAS) in COUNSELINGEVAL assess empathy and coherence (Lee et al., 2024a), and interactional features like PQA are explored (Shen et al., 2024). However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW), which are potentially useful for structured PE flows, remain underused. While some qualitative reviews consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and structured questioning is explored (Ren et al., 2024), robust frameworks for quantitative evaluation of PE-specific structure are lacking.
@@ 3.4 Bridging the Gap in PE Therapy Evaluation
PE is a first-line treatment for PTSD (Sherrill and Rauch, 2019; Varkovitzky et al., 2018; Rauch et al., 2021; Ragsdale et al., 2020; Evans et al., 2020; Yasinski et al., 2017). Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation, a gap remains in fidelity assessment for trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024). Current approaches lack metrics and focus tailored to PE’s unique conversational structure, including avoidance handling, trauma cue processing, and imaginal exposure flow, and they do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024). Our work systematically evaluates synthetic data for alignment with PE structure and clinical validity, identifying areas for improving synthetic PE dialogue quality (see Table 1, Appendix A).
@@ 4 Methodology
We analyzed 400 prolonged exposure therapy conversations, consisting of 200 real-world and 200 synthetic. To demonstrate our evaluation framework, this case study uses 200 synthetic conversations from the publicly available Thousand Voices of Trauma dataset (BN et al., 2025b). The data was generated using Claude Sonnet 3.5 (Anthropic, 2024) with PE-specific prompting frameworks adapted from clinical guidelines. This dataset includes diverse therapist-client interactions across trauma types, therapy phases, and demographics.
The real-world PE sessions were collected under IRB-approved protocols, with participant consent, and cannot be released due to privacy constraints. Each session (1-1.5 hours) was transcribed using Amazon HealthScribe (Services, 2023), manually verified, and reformatted by merging consecutive speaker turns for readability. Each conversation followed a validated therapy fidelity checklist (See Table 1) to align with real-world standards. Both datasets underwent the same preprocessing:
@@ 4.1 Metrics and Analysis
To systematically compare real and synthetic therapeutic conversations, we selected a diverse set of linguistic, structural, and statistical metrics. These metrics provide insights into conversational dynamics, protocol adherence, and overall fidelity, ensuring a holistic evaluation of synthetic dialogue generation. To reduce subjectivity, we adopt a fidelity checklist (Table 1) and PE-specific metrics grounded in existing clinical guidelines for evaluating therapeutic adherence. Our methodology consists of four key analyses: system-level metrics comparison, correlation analysis, statistical significance testing, and PE-specific metrics.
@@ 4.1.2 Correlation Analysis
While mean comparisons provide a general overview, correlation analysis (see Table 2) quantifies the consistency of relationships across metrics between real and synthetic data.
@@ 4.1.3 Statistical Significance Testing
To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 3).
To identify which properties most strongly differentiate real and synthetic sessions, we next examine feature importance from a simple predictive model (see Table 4).
@@ 4.1.5 Qualitative Fidelity Assessment
To supplement automated evaluation metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using established PE clinical guidelines. Manual fidelity annotations were then reviewed by a licensed clinical psychotherapist for adherence to PE protocol. While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses, especially those potentially overlooked by automated metrics, rather than as quantitative ground truth. Representative examples of such lapses identified through this process are illustrated in Figure 1 and discussed in Section 6.2.
@@ 5.1 System-Level Metrics & Correlation
    Concise and structured responses: Synthetic dialogues are shorter (Average Utterance Length: 22.90 $\pm$ 1.74 vs. 68.72 $\pm$ 26.61) and more consistent (Utterance Length Std Dev: 18.54 $\pm$ 2.35 vs. 135.85 $\pm$ 66.25) due to LLM output constraints. Techniques like Chain-of-Thought prompting help improve coherence, though larger output contexts are needed.
@@
    Correlation analysis: High correlations in turn-taking metrics (0.65-0.78 synthetic vs. 0.85-0.92 real) confirm accurate conversational dynamics. Lower correlations in linguistic complexity metrics (e.g., Readability Score, Vocabulary Richness, Perplexity) reflect structural differences but do not hinder the coherence/naturalness of the conversation.
@@
Norm. Total Turns $\pm 0.0 1.00±0.01$$\pm 0.0$ $0.78$ $0.65$
@@
Turn Ratio (T/C) $01\pm 0.0 0.0$$01\pm 0.0$ $0.80$ $0.66$
Word Ratio (T/C) $1\pm 0.0 0.00±0$$0\pm 0.0$ $0.82$ $0.68$
@@ 5.2 Statistical Significance Testing
For example, while a real client might repeatedly use an informal phrase like “freaked out,” the LLM generates diverse but semantically similar alternatives such as “felt anxious,” “was overcome with fear,” or “grew very distressed,” thereby increasing lexical richness within a tight, clinically-appropriate boundary.
@@
T-C Turn Ratio $3\times 10^{3} 1.22×1$$2\times 10^{-9}$
@@ 5.3 Feature Importance
Feature importance analysis (see Table 4) identifies the metrics most influencing synthetic data generation, confirming trends from previous evaluations. While not the central focus, these results support the model’s alignment with conversational structure and linguistic patterns.
@@
    Lexical richness is moderately important: Synthetic dialogues use varied yet structured language, balancing vocabulary diversity with clarity, while real conversations are more flexible.
@@ 5.4 Prolonged Exposure Specific Metrics
Table 5 presents the statistical test results comparing synthetic Prolonged Exposure data with real therapy session data. The synthetic data shows promise, with no statistically significant difference from real sessions on key temporal metrics like Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$). Similarly, Narrative Development ($p=0.251$) appears less robust but also shows no significant statistical difference, possibly due to limited context retention. This suggests the model is beginning to capture the therapeutic arc of distress reduction. Conversely, foundational constructs such as Trauma Narrative Coherence, Emotional Engagement, and Avoidance Handling show statistically significant differences (p < 0.001) when compared to real dialogues. These findings provide a clear and actionable roadmap for future work, guiding efforts to calibrate the model’s consistency with the variability of real-world therapeutic interactions. Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications. Addressing dynamic narrative evolution could further improve alignment with real data. However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics. For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring. These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.